The health of a lion population in a safari park is monitored by measurements. a sample of 86 lions has a mean weight of 169.8 lb. Assume that the population standard deviation is known to be 46.5 lb and find a 93% confidence interval estimate of the mean of the population and find the margin of error.
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 169.8

Population standard deviation =    = 46.5

Sample size = n = 86

At 93% confidence level

= 1 - 93%  

= 1 - 0.93 =0.07

/2 = 0.035

Z/2 = Z0.035 = 1.812


Margin of error = E = Z/2 * ( /n)

= 1.812 * ( 46.5 /  86 )

= 9.09

At 93% confidence interval estimate of the population mean is,

  ± E

169.8 ± 9.09

(160.71, 178.89)  

Margin of error = E = Z/2 * ( /n)

E = 1.812 * ( 46.5 /  86 )

E = 9.09